What volume of gaseous chlorine at STP would you need to obtain a 269.5-g sample of gas?
1
Expert's answer
2020-11-17T10:42:02-0500

M(Cl2) = 70.9 g/mol

"n = \frac{m}{M}"

n(Cl2) "= \frac{269.5}{70.9} = 3.8 \;mol"

1 mole of any gas occupies 22.4 L at STP.

V(Cl2) "= 3.8 \times 22.4 = 85.12\; L"